the locus of the point (4t^2 -1 , 8t-2 ) represents a parabola then the equation of latus rectum is???
please give solution

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
x = 4t² -1
y = 8t -2
 
x +1 = 4 (y+2/8)²
 
16 x +16 = y² +4 + 4y
(y +2)² = 16 (x+1)
 
Hence
 
Length of latus rectum = 16
Also equation is x +1 = 4
x = 3